I'm dealt 7 cards from a standard deck of 52 cards.
In how many different ways can all cards be of the same suit?

Answer by stanbon(75887)

I'm dealt 7 cards from a standard deck of 52 cards.
In how many different ways can all cards be of the same suit?
------------------------
Ans: 4*13C7 = 4*1716 = 6864
---------------------------------
Cheers,
Stan H.
--------------